OpenStudy (anonymous):

EASY MEDAL/FAN Solve for x: −3|2x + 6| = −12 A)x = 1 and x = 5 B)x = −1 and x = −5 C)x = −9 and x = 3 D)No solution

OpenStudy (anonymous):

69+69 = d i c k up ur a s s c u n t SLAM F U C K I N G N I G G E R S INTO THE GROUND!

OpenStudy (decentnabeel):

\[-3\left|2x+6\right|=-12\quad :\quad x=-5\:\:\:\mathrm{or}\:\:\:\:x=-1\]

OpenStudy (decentnabeel):

option B is the answer

OpenStudy (decentnabeel):

Are you understand @shawty247

OpenStudy (anonymous):

YEA THANKS

OpenStudy (decentnabeel):

no problem